THE RENNER-TELLER AND JAHN-TELLER EFFECTS IN PROTOTYPICAL MOLECULAR CATIONS SUBJECT TO A VERY LARGE SPIN-ORBIT COUPLING

\underline{BÉRENGER GANS}, MONIKA GRÜTTER and FRÉDÉRIC MERKT, Laboratorium für Physikalische Chemie, ETH Zürich, CH-8093 Zürich, Switzerland.

PFI-ZEKE photoelectron spectra of the X^{+} ^{2}Π ← X ^{1}Σ^{+} transition of HC_{2}I and of the X̃^{+} ^{2}E_{3/2} ← X̃ ^{1}A_{1} transition of CH_{3}I have been recorded at a resolution sufficiently high to observe, at least partially, their rotational structure. The spin-rovibronic energy-level structures of HC_{2}I^{+} and CH_{3}I^{+} could be determined at low energies and enabled us to study the Renner-Teller and Jahn-Teller effects in molecular cations subject to a very large spin-orbit coupling with unprecedented details. In the case of HC_{2}I, the nominally forbidden 5_{0}^{1} band has been observed in addition to the origin band and allowed us to determine a splitting of 2 cm^{−1} between the two Renner-Teller components of the 5^{1} vibrational level of the cation. In the case of CH_{3}I, the rotational structure of the origin and of the 2_{0}^{1} and 3_{0}^{1} bands are dominated by satellite bands of spin-rovibronic origin^{a}. The 5_{0}^{1} and 6_{0}^{1} bands reveal an additional splitting corresponding to the separation between the two Jahn-Teller components of j = 1/2 and j = 3/2 symmetry of the 5^{1} and 6^{1} levels of the cation^{b}.

aM. Grütter, J.M. Michaud and F. Merkt, J. Chem. Phys. 134, 054308 (2011).

bT.A. Barckholtz and T.A. Miller, Int. Rev. Phys. Chem. 17 (4), 435 (1998).